# Robust Universal Inference For Misspecified Models

### v3

@@ Abstract
In statistical inference, it is rarely realistic that the hypothesized statistical model is well-specified, and consequently, it is important to understand the effects of misspecification on inferential procedures. When the hypothesized statistical model is misspecified, the natural target of inference is a projection of the data generating distribution onto the model. We present a general method for constructing valid confidence sets for such projections, under weak regularity conditions, despite possible model misspecification. Our method builds upon the universal inference method and is based on inverting a family of split-sample tests of relative fit. We study settings in which our methods yield either exact or approximate, finite-sample valid confidence sets for various projection distributions. We study rates at which the resulting confidence sets shrink around the target of inference and complement these results with a simulation study and a causal discovery using linear causal model on CausalEffectPairs dataset.
@@ 1 Introduction
In parametric models minimum distance methods [10, 6] explicitly aim to estimate projection parameter $\widetilde{\theta}$. The maximum likelihood estimator can be viewed as a minimum distance estimator when taking $\rho$ to be the Kullback-Leibler (KL) divergence. However, absent strong regularity conditions, the KL projection is highly non-robust to small perturbations in $P^{*}$, even tiny perturbations to $P^{*}$ can lead to vastly different projections $\widetilde{P}$ [10, 6]. From a practical standpoint, these instabilities can make the KL projection an undesirable target, and in these cases it is essential to develop a flexible family of methods that can target other (more stable) projection distributions. Each choice of the divergence yields a different target and in most cases, these targets will have drastically different properties. This in turn poses significant challenges in constructing a unified framework for statistical inference in the misspecified setting.
@@
[…]
@@
In recent work [58], we introduced a universal inference procedure (described in more detail in Section 2) based on data-splitting to construct uniformly, finite-sample valid likelihood-ratio confidence sets under no regularity conditions. This work showed that, in the well-specified setting, sample-splitting can yield practical, finite-sample valid inference, even for irregular statistical models, often at a surprisingly small statistical price. As highlighted by recent work [54, 23, 52] sample-splitting reduces efficiency compared to standard methods in regular problems and are often conservative in these settings. Universal inference, and the methods we propose in this paper, are most useful for irregular problems where standard methods fail.
@@ 2.2 A Re-Interpretation of Universal Inference
for a choice of the threshold $t_{\alpha}(P,\widehat{P}_{1})$ which ensures that the condition in (4) is satisfied. Although it is possible to determine optimal thresholds in the likelihood-ratio tests above, this can be practically cumbersome since these thresholds depend on both the pilot estimate $\widehat{P}_{1}$ and the null hypothesis $P$ under consideration. The work of Wasserman et al. [58] further shows that a universal threshold $t_{\alpha}=1/\alpha$ suffices to ensure the condition in (4). To summarize, one can view the universal inference confidence set (3) as arising by inverting a family of likelihood-ratio tests designed to distinguish each candidate distribution $P$ from a pilot estimate $\widehat{P}_{1}$.
@@ 2.3 Divergences
[…]
@@
[…]
and $\mathsf{DP}_{0}=\mathsf{KL}$ is defined by taking the limit of $\beta\to 0$. When $\beta=1$, the density power divergence is the $L_{2}$ distance $\int(p-q)^{2}\,{\mathrm{d}}\lambda$. Finally, the family of Integral Probability Metrics [IPMs, 45], are defined as
@@
where $\mathcal{F}$ is a symmetric class (i.e., $f\in\mathcal{F}\iff-f\in\mathcal{F}$) of real-valued bounded measurable functions on the domain of $P$ and $Q$. Important special cases of IPMs include the Total Variation distance (TV, where $\mathcal{F}$ is the collection of functions with sup-norm at most 1), the Wasserstein-1 distance (where $\mathcal{F}$ is the collection of 1-Lipschitz functions) and the Maximum Mean Discrepancy (where $\mathcal{F}$ is the unit ball of a Reproducing Kernel Hilbert Space).
@@ 3 Model Misspecification and Unstable Projection Parameters
To provide some motivation and intuition for the methods we propose in this paper, it is useful to understand some of the failures of the universal inference framework when the statistical model is misspecified, and the target of inference is the KL projection. All proofs of the examples given in this section are in Supplementary Material B.
@@ 3.1 Unbounded Likelihood-Ratios
[…]
@@
[…]
@@
Huber’s $\epsilon$-contamination model yields a canonical illustration of the non-robustness of the KL projection. The following example shows that universal inference can also fail in this setting.
@@
Suppose we observe $X_{1},\ldots,X_{n}\sim P^{*}:=(1-\epsilon_{n})N(0,1)+\epsilon_{n}Q_{a_{n}}$ for some small positive $0<\epsilon_{n}\leq c_{1}/n$, $|a_{n}|\geq c_{2}n^{2}$ where $c_{1},c_{2}>0$ are small positive constants, and $Q_{a}$ is uniform on $[a-\delta,a+\delta]$ for some $\delta>0$. Consider a Gaussian location model $\mathcal{P}=\left\{N(\theta,1):\theta\in\mathbb{R}\right\}$. Suppose we take the pilot estimator to be the MLE (6). Then, for all sufficiently large $n\geq n_{0}$ where $n_{0}$ only depends on $\alpha$ the split LRT confidence set in (3), with an equal sized split into $\mathcal{D}_{0}$ and $\mathcal{D}_{1}$, fails to cover the KL projection $\widetilde{P}$ at the nominal level.
[…]
@@ 3.2 Failure Despite Bounded Likelihood-Ratios
[…]
@@ 4.3 Approximate Robust Universal Confidence Sets
The set $\widetilde{P}_{\nu}$ is a collection of distributions that are as close to $P^{*}$ as $\widetilde{P}$ (up to a factor $\nu$). The approximate confidence set guarantee is most meaningful when $\nu$ is close to 1, or when the model misspecification is not too extreme, i.e. $\rho(P^{*}\|\widetilde{P})$ is small. The approximate confidence set requires some care to interpret. In practice we are sometimes interested in constructing a confidence set for a functional $\psi:\mathcal{P}\mapsto\mathbb{R}^{d}$ of the projection distribution (say, its mean). A $\nu$-approximate confidence set for $\widetilde{P}$ does not in general yield an approximate confidence set for $\psi(\widetilde{P})$ except under strong regularity conditions which allow us to relate the divergence $\rho$ between two distributions to a distance between their induced functional values (for instance, differentiability in quadratic mean [56]).
@@
The critical shortcoming of a $\nu$-approximate confidence set is that its interpretation is challenging as it does not guarantee coverage of the projection distribution $\widetilde{P}$. In some cases, for example, when the model misspecification is a consequence of data contamination, the statistician might be able to provide a reasonable upper bound on $\rho(P^{*}\|\widetilde{P})$. As an example when the distribution $P^{*}$ arises from an $\epsilon$-contamination of a distribution $P\in\mathcal{P}$ then we have the upper bound $\text{TV}(P^{*},\widetilde{P})\leq\epsilon$. In this case one can construct a valid (classical) confidence set by enlarging the $\nu$-approximate confidence set. For a given upper bound $\epsilon$ on $\rho(P^{*}\|\widetilde{P})$ define:
@@ 5.1 Kullback-Leibler Divergence
where $\mathcal{I}_{0}$ is an index set of $\mathcal{D}_{0}$ and $t_{\alpha}(P,\widehat{P}_{1})$ is chosen to ensure validity. This test was called the relative information fit test (RIFT) and studied in the work of Chakravarti et al. [18] to study the relative goodness-of-fit of two candidate estimates. In our paper, we invert the same test in order to construct a robust universal confidence set.
@@
$\sqrt{n_{0}}\left(\overline{T}_{n_{0}}(P,\widehat{P}_{1})-\mathbb{E}_{P^{*}}T(P,\widehat{P}_{1})\right)\overset{P^{*}}{\rightsquigarrow}\mbox{$\mathsf{N}$}\left(0,s_{P}^{2}\right)\qquad\text{ as }n_{0}\to\infty$
@@ 5.2 Density Power (DP) Divergences
$\overline{T}_{n_{0}}(P,\widehat{P}_{1})$ =$=\int\left\{p^{1+\beta}-\widehat{p}_{1}^{1+\beta}\right\}\,{\mathrm{d}}\lambda-\left(1+\frac{1}{\beta}\right)\frac{1}{n_{0}}\sum_{i\in\mathcal{I}_{0}}\left[p^{\beta}-\widehat{p}_{1}^{\beta}\right](X_{i}).$ (21)
@@
[…]
@@ 5.3 Hellinger Distance
Our strategy will be to modify the $\rho$-estimator [^2] The name “$\rho$-estimator” comes from the standard symbol used for the Hellinger affinity. [3, 2, 5] which is a density estimator tailored to the Hellinger loss. Define the split $\rho$-test statistic
@@ 6 Size of Robust Universal Confidence Sets
Theorem 5 also shows how the size of the set depends on the choice of $\widehat{P}_{1}$. When possible we should choose a pilot estimate $\widehat{P}_{1}$ which converges to the target $\widetilde{P}$ at a fast rate to ensure that the term $\eta_{n}$ is sufficiently small. A sensible choice is often a minimum distance estimator [MDE, 46] which is not only a consistent estimator of $\widetilde{P}$ under some regularity conditions but is also robust to some misspecification in its corresponding distance [43, 22].
@@ 7 Practical Implementation
For most parametric models, $pv(\theta)$ is infinitely differentiable so that $\gamma=\infty$ and hence as long as $B\log n/n\to\infty$, we have that $\mathbb{P}_{P^{*}}(\widetilde{\theta}(P)\in C_{\alpha})=1-\alpha+O_{P^{*}}(n^{-1/2}).$
@@ 8 Illustration: Contaminated Data Generating Process
Figure 4 illustrates the empirical coverage and size of split LRT and ($\mathsf{KL}$ and $\mathsf{DP}$) ReDI sets based on 1000 replications. For split LRT and KL ReDI sets, we choose $\widehat{\theta}_{1}$ to be the quasi-MLE of the first half samples, whereas, for the DP ReDI set, we use the minimum DP divergence estimator. The non-robust split LRT fails to cover KL projection in all cases whereas ReDI sets achieve the nominal coverage with large enough sample sizes. The DP ReDI sets show superior coverage than KL ReDI set across all sample sizes. Such a target coverage improvement is more evident in the smaller sample sizes below 200, and as $\beta$ gets larger, i.e., the DP ReDI set targets a more stable projection. Regardless of what divergence measure $\rho$ is of interest, the size of the confidence set with reference to $\rho$ shrinks to zero as the sample size increases. Again, the extremal values of $\mathsf{KL}_{\mathscr{H}}^{P^{*}}(\widehat{C}_{\textsc{ReDI}},\widetilde{P})$ for sample sizes below 500 highlight the instability of KL projection.
@@ 9 Application: Causal Discovery with Homoscedastic Linear Causal Model
We used the CausalEffectPairs [44] dataset, which offers a collection of cause-effect pairs from diverse domains, with ground-truth causal directions established through domain expertise. For our application, we consider three representative pairs: pair 76 (the average annual growth of population $X_{1}$ and the food consumption $X_{2}$), pair 18 (age of child $X_{1}$ and chemical GAG concentration in urine $X_{2}$), and pair 8 (abalone age $X_{1}$ and height $X_{2}$). These examples were selected to illustrate varying degrees and types of model misspecification. Pair 76 is considered nearly well-specified and was also analyzed in Strieder et al. [53]. Pair 18 exhibits nonlinearity and possible deviations from homoscedasticity, while pair 8 involves both nonlinearity and heteroscedasticity, as well as notable outliers. See Figure 6 for scatterplots of each pairs after centering.
@@
Figure 7 shows the 95% confidence sets for the total causal effects $C(1\to 2)$ and $C(2\to 1)$, constructed using two LRT variants [LRT1 and LRT2, 53], split LRT, KLReDI, DPReDI with $\beta=0.1,1$. For split LRT and ReDI sets, we provide both vanilla variant—based on equally split data—and the crossfit variant, which averages the split test statistics and its corrupted variances by swapping the role of split samples $\mathcal{D}_{0}$ and $\mathcal{D}_{1}$. This crossfitting approach can help reduce split-sample variability and improve efficiency. See Supplementary Material G for details. All methods correctly identified the ground-truth causal direction across all three pairs. LRT2 fails to provide confidence set in all cases, despite its strong simulation performance, whereas LRT1 yielded the narrowest confidence intervals.
@@
For pair 8, where the data exhibit nonlinearity, heteroscedasticity, and contain outliers, KLReDI produces significantly wider confidence sets than split LRT, reflecting the impact of model misspecification. In contrast, the $\mathsf{DP}\textsc{ReDI}$ set is comparable in width to, or even narrower than, split LRT. This highlights the sensitivity of KL projection to small but adversarial perturbations (as discussed in Example 3 and Section 8), and underscores the benefit of targeting a more stable projection for inference. The influence of the DP parameter $\beta$ is especially clear in this example: as $\beta$ increases, the confidence set for $C(2\to 1)$ shifts toward larger values. This is due to the increased robustness of the DP projection, which more effectively downweights outliers (e.g., $(X_{1},X_{2})=(-1.9,1.0)$). Finally, note that the LRT1 and cross-fit $\mathsf{DP}\textsc{ReDI}_{1}$ sets do not overlap, indicating that inference based on an overconfident LRT—without accounting for model misspecification—can substantially underestimate the total causal effect $C(2\to 1)$.

### v4

@@ Abstract
In statistical inference, it is rarely realistic that the hypothesized statistical model is well-specified, and consequently, it is important to understand the effects of misspecification on inferential procedures. When the hypothesized statistical model is misspecified, the natural target of inference is a projection of the data generating distribution onto the model. We present a general method for constructing valid confidence sets for such projections, under weak regularity conditions, despite possible model misspecification. Our method builds upon the universal inference method and is based on inverting a family of split-sample tests of relative fit. We study settings in which our methods yield either exact or approximate, finite-sample valid confidence sets for various projection distributions. We study rates at which the resulting confidence sets shrink around their target of inference and complement these results with a simulation study and a study of causal discovery using a linear causal model with the CausalEffectPairs dataset.
@@ 1 Introduction
In parametric models minimum distance methods [10, 6] explicitly aim to estimate projection parameter $\widetilde{\theta}$. The maximum likelihood estimator can be viewed as a minimum distance estimator when taking $\rho$ to be the Kullback-Leibler (KL) divergence. However, absent strong regularity conditions, the KL projection is highly non-robust to small perturbations in $P^{*}$, even tiny perturbations to $P^{*}$ can lead to vastly different projections $\widetilde{P}$ [10, 6]. From a practical standpoint, these instabilities can make the KL projection an undesirable target, and in these cases it is essential to develop a flexible family of methods that can target other (more stable) projection distributions. Each choice of the divergence yields a different target. In some cases, these targets will have drastically different properties. This in turn poses significant challenges in constructing a unified framework for statistical inference in the misspecified setting.
@@
Suppose we observe $Y_{1},\ldots,Y_{n}\overset{iid}{\sim}P^{*}$ where $Y_{i}\in\{0,1,\ldots\}$, and $P^{*}$ is a Negative Binomial distribution with mean 4 and variance 18. We want to approximate $P^{*}$ with a mixture of Geometric distributions $(1-\pi)\mbox{$\mathsf{Geom}$}(\lambda_{1})+\pi\mbox{$\mathsf{Geom}$}(\lambda_{2})$ where $\pi\in[0,1]$ and $\mbox{$\mathsf{Geom}$}(\lambda)$ denotes a Geometric with mean $\lambda\in(0,1]$. The parameter is $\theta=(\pi,\lambda_{1},\lambda_{2})$. This model fails to satisfy the usual regularity conditions and $\theta$ is not identifiable in general. Our choice of $P^{*}$ is well-approximated by a single Geometric distribution, and consequently the $L^{2}$ projection parameter $\widetilde{\theta}$ is a set of parameter values $\widetilde{\theta}_{L_{2}}=([0,1],0.19,0.19)$. Standard asymptotic methods for constructing confidence sets for the projection parameter fail. Figure 1 shows a valid confidence set for $L_{2}$ projection parameter $\widetilde{\theta}$ computed using the numerical algorithm in Section 7. This unusual confidence set correctly mitigates the two main issues — the lack of identifiability and model misspecification — to yield a valid confidence set. Specifically, the confidence set shows that if $\lambda_{1}\approx 0.2$ then $\lambda_{2}$ is essentially unconstrained and vice versa.
@@
In recent work [59], we introduced a universal inference procedure (described in more detail in Section 2) based on data-splitting to construct uniformly, finite-sample valid likelihood-ratio confidence sets under no regularity conditions. This work showed that, in the well-specified setting, sample-splitting can yield practical, finite-sample valid inference, even for irregular statistical models, often at a surprisingly small statistical price. As highlighted by recent work [55, 24, 53] sample-splitting reduces efficiency compared to standard methods in regular problems and universal inference is often conservative in these settings. Universal inference, and the methods we propose in this paper, are most useful for irregular problems where standard methods fail.
@@ 2.2 A Re-Interpretation of Universal Inference
for a choice of the threshold $t_{\alpha}(P,\widehat{P}_{1})$ which ensures that the condition in (4) is satisfied. Although it is possible in principle to determine optimal thresholds in the likelihood-ratio tests above, this can be practically cumbersome since these thresholds depend on both the pilot estimate $\widehat{P}_{1}$ and the null hypothesis $P$ under consideration. The work of Wasserman et al. [59] further shows that a universal threshold $t_{\alpha}=1/\alpha$ suffices to ensure the condition in (4). To summarize, one can view the universal inference confidence set (3) as arising by inverting a family of likelihood-ratio tests designed to distinguish each candidate distribution $P$ from a pilot estimate $\widehat{P}_{1}$.
@@ 2.3 Divergences
The Hellinger distance is defined as: ${\sf H}(P,Q)=\frac{1}{\sqrt{2}}\left(\int(\sqrt{p}-\sqrt{q})^{2}d\lambda\right)^{1/2},$ and the Kullback-Leibler (KL) divergence is defined as:
@@
$\mathsf{DP}_{\beta}(P\|Q)$ $=\int\left\{q^{1+\beta}-\left(1+\frac{1}{\beta}\right)q^{\beta}p+\frac{1}{\beta}p^{1+\beta}\right\}d\lambda,$ $\beta>0$
and $\mathsf{DP}_{0}=\mathsf{KL}$ is defined by taking the limit of $\beta\to 0$. When $\beta=1$, the density power divergence is the $L_{2}$ distance $\int(p-q)^{2}\,{\mathrm{d}}\lambda$. Finally, the family of Integral Probability Metrics [IPMs, 46], are defined as
@@
where $\mathcal{F}$ is a symmetric class (i.e., $f\in\mathcal{F}\implies-f\in\mathcal{F}$) of real-valued bounded measurable functions on the domain of $P$ and $Q$. Important special cases of IPMs include the Total Variation distance (TV, where $\mathcal{F}$ is the collection of functions with sup-norm at most 1), the Wasserstein-1 distance (where $\mathcal{F}$ is the collection of 1-Lipschitz functions) and the Maximum Mean Discrepancy (where $\mathcal{F}$ is the unit ball of a Reproducing Kernel Hilbert Space).
@@ 3 Model Misspecification and Unstable Projection Parameters
To provide some motivation and intuition for the methods we propose in this paper, it is useful to understand some of the failures of the universal inference framework when the statistical model is misspecified, and the target of inference is the KL projection. All proofs accompanying the examples given in this section are in Supplementary Material B.
@@ 3.1 Unbounded Likelihood-Ratios
The behavior of likelihood-ratio based methods can be sensitive to the tail behavior of likelihood-ratios. The following simple example illustrates that under model misspecification, universal inference can fail to cover the KL projection parameter. These pathologies do not arise when the statistical model is correctly specified, and the challenges in this example arise due to an interplay between poorly behaved likelihood-ratios and model misspecification. This example also serves to highlight the fact that the KL projection parameter can in some cases be an undesirable inferential target. We let $\mbox{$\mathsf{Bern}$}(p)$ denote the Bernoulli distribution with parameter $p$.
@@
In this example the KL projection distribution $\widetilde{P}$ is $\mbox{$\mathsf{Bern}$}(1/2)$. For $\epsilon_{n}\ll 1/n$, with high probability the samples $X_{1},\ldots,X_{n}$ are all 0. Consequently, the MLE $\widehat{P}_{1}$ with high-probability will be $\mbox{$\mathsf{Bern}$}(0)$. Furthermore, the split sample likelihood $\mathcal{L}_{0}$ will be much higher for $\mbox{$\mathsf{Bern}$}(0)$ than $\mbox{$\mathsf{Bern}$}(1/2)$, and consequently $\mbox{$\mathsf{Bern}$}(1/2)$ will not be included in the universal set.
@@
Huber’s $\epsilon$-contamination model yields another canonical illustration of the non-robustness of the KL projection. The following example shows that universal inference can fail in this setting.
@@
Suppose we observe $X_{1},\ldots,X_{n}\sim P^{*}:=(1-\epsilon_{n})N(0,1)+\epsilon_{n}Q_{a_{n}}$ for some small positive $0<\epsilon_{n}\leq c_{1}/n$, $|a_{n}|\geq c_{2}n^{2}$ where $c_{1},c_{2}>0$ are small positive constants, and $Q_{a}$ is uniform on $a-\delta,a+\delta]$ for some $\delta>0$. Consider a Gaussian location model $\mathcal{P}=\left\{N(\theta,1):\theta\in\mathbb{R}\right\}$. Suppose we take the pilot estimator to be the MLE (6). Then, for all sufficiently large $n\geq n_{0}$ where $n_{0}$ only depends on $\alpha$ the split LRT confidence set in (3), with an equal sized split into $\mathcal{D}_{0}$ and $\mathcal{D}_{1}$, fails to cover the KL projection $\widetilde{P}$ at the nominal level.
The KL projection parameter in this example is $\widetilde{\theta}=\epsilon_{n}a_{n}$ which tends to $\infty$ as $n\to\infty$ despite $\epsilon_{n}\ll 1/n$. With high probability, all samples are nearly indistinguishable from those of $\mbox{$\mathsf{N}$}(0,1)$. The MLE $\widehat{\theta}_{1}=\overline{X}_{n_{1}}$ will be much closer to zero than $\widetilde{\theta}$, resulting in a higher split sample likelihood $\mathcal{L}_{0}$ for $\widehat{\theta}_{1}$ than $\widetilde{\theta}$. Consequently $\widetilde{\theta}$ will not be included in the universal set. The fundamental issue again lies in the instability of the KL projection, which arises from the sensitivity of the expected log-likelihood ratio to its tail behavior. In contrast to the KL projection, the Hellinger projection, for example, tends to the uncontaminated distribution $N(0,1)$ [10] as $\epsilon_{n}\rightarrow 0$ and is a better inferential target. In Section 8, we empirically demonstrate this issue and show that targeting a more stable projection alleviates the problem.
@@ 3.2 Failure Despite Bounded Likelihood-Ratios
The KL projection distribution is $\mbox{$\mathsf{Bern}$}(3/4)$. We show that the pilot estimate $\widehat{P}_{1}$ with probability near $1/2$ will be the distribution $\mbox{$\mathsf{Bern}$}(1/4)$, and further with probability near $1/2$ the KL projection $\mbox{$\mathsf{Bern}$}(3/4)$ will have a much smaller split sample likelihood than $\widehat{P}_{1}$. As a direct consequence, universal inference will fail to cover the projection distribution $\mbox{$\mathsf{Bern}$}(3/4)$.
@@ 4.3 Approximate Robust Universal Confidence Sets
[…]
@@
[…]
@@ 5.1 Kullback-Leibler Divergence
where $\mathcal{I}_{0}$ is an index set of $\mathcal{D}_{0}$ and $t_{\alpha}(P,\widehat{P}_{1})$ is chosen to ensure validity. This test was called the relative information fit test (RIFT) and was used to select between two candidate estimates by Chakravarti et al. [19]. In our paper, we invert the same test in order to construct a robust universal confidence set.
@@
[…]
@@ 5.2 Density Power (DP) Divergences
$\overline{T}_{n_{0}}(P,\widehat{P}_{1})$ $=\int\left\{p^{1+\beta}-\widehat{p}_{1}^{1+\beta}\right\}\,{\mathrm{d}}\lambda-\left(1+\frac{1}{\beta}\right)\frac{1}{n_{0}}\sum_{i\in\mathcal{I}_{0}}\left[p^{\beta}-\widehat{p}_{1}^{\beta}\right](X_{i}).$ (21)
@@
A formal proof can be found in Supplementary Material D. The key observation is that the DP projection is $\mbox{$\mathsf{Bern}$}(0)$ for a sufficiently large sample size for any fixed $\beta>0$. The DP projection in this example is more stable than the KL projection $\mbox{$\mathsf{Bern}$}(1/2)$, considering that $\epsilon_{n}$ is much closer to 0 than $1/2$. Consequently, we show that the DP ReDI set will cover the target of inference $\mbox{$\mathsf{Bern}$}(0)$ with high probability. We emphasize that the MLE is also $\mbox{$\mathsf{Bern}$}(0)$ with high probability, yet both universal split LRT and KL ReDI set based on the MLE fail to cover the KL projection due to the instability of the population projection distribution.
@@ 5.3 Hellinger Distance
Our strategy will be to modify the $\rho$-estimator [^2] The name $\rho$-estimator” comes from the standard symbol used for the Hellinger affinity. [3, 2, 5] which is a density estimator tailored to the Hellinger loss. Define the split $\rho$-test statistic
@@ 6 Size of Robust Universal Confidence Sets
Theorem 5 also shows how the size of the set depends on the choice of $\widehat{P}_{1}$. When possible we should choose a pilot estimate $\widehat{P}_{1}$ which converges to the target $\widetilde{P}$ at a fast rate to ensure that the term $\eta_{n}$ is sufficiently small. A sensible choice is often a minimum distance estimator [MDE, 47] which is not only a consistent estimator of $\widetilde{P}$ under some regularity conditions but is also robust to some misspecification in its corresponding distance [44, 23].
@@ 8 Illustration: Contaminated Data Generating Process
Figure 4 illustrates the empirical coverage and size of split LRT and ($\mathsf{KL}$ and $\mathsf{DP}$) ReDI sets based on 1000 replications. For split LRT and KL ReDI sets, we choose $\widehat{\theta}_{1}$ to be the quasi-MLE of the first half samples, whereas, for the DP ReDI set, we use the minimum DP divergence estimator. The non-robust split LRT fails to cover KL projection in all cases whereas ReDI sets achieve the nominal coverage with large enough sample sizes. The DP ReDI sets show superior coverage than KL ReDI set across all sample sizes. Such a target coverage improvement is more evident in the smaller sample sizes below 200, and as $\beta$ gets larger, i.e., the DP ReDI set targets a more stable projection. Regardless of what divergence measure $\rho$ is of interest, the size of the confidence set with reference to $\rho$ shrinks to zero as the sample size increases. Again, the extremal values $\mathsf{KL}_{\mathscr{H}}^{P^{*}}(\widehat{C}_{\textsc{ReDI}},\widetilde{P}) fo$r sample sizes below 500 highlight the instability of KL projection.
@@ 9 Application: Causal Discovery with Homoscedastic Linear Causal Model
We used the CausalEffectPairs [45] dataset, which offers a collection of cause-effect pairs from diverse domains, with ground-truth causal directions established through domain expertise. For our application, we consider three representative pairs: pair 76 (the average annual growth of population $X_{1}$ and the food consumption $X_{2}$), pair 18 (age of child $X_{1}$ and chemical GAG concentration in urine $X_{2}$), and pair 12 (age $X_{1}$ and wage per hour $X_{2}$). These examples were selected to illustrate varying degrees and types of model misspecification. Pair 76 is considered nearly well-specified and was also analyzed in Strieder et al. [54]. Pair 18 exhibits nonlinearity and possible deviations from homoscedasticity, while pair 12 involves both nonlinearity and heteroscedasticity, as well as notable outliers. See Figure 6 for scatterplots of each pairs after centering.
@@
Figure 7 shows the 95% confidence sets for the total causal effects $C(1\to 2)$ and $C(2\to 1)$, constructed using two LRT variants [LRT1 and LRT2, 54], split LRT, KLReDI, DPReDI with $\beta=0.1,1$. For split LRT and ReDI sets, we provide both vanilla variant—based on equally split data—and the crossfit variant, which averages the split test statistics and its corrupted variances by swapping the role of split samples $\mathcal{D}_{0}$ and $\mathcal{D}_{1}$. This crossfitting approach can help reduce split-sample variability and improve efficiency. See Supplementary Material G for details. All methods correctly identified the ground-truth causal direction across all three pairs. LRT2 fails to provide confidence set in all cases, despite its strong simulation performance, whereas LRT1 yielded the narrowest confidence intervals.
@@
For pair 12, where the data exhibit nonlinearity, heteroscedasticity, and contain outliers, the $\mathsf{KL}$ReDI produced a wider confidence set than the split LRT as in pair 18. The $\mathsf{DP}\textsc{ReDI}_{0.1}$ set was narrower than both split LRT and $\mathsf{KL}$ReDI sets, while remaining similarly centered, indicating a reduced variance from a small robustness gain. In contrast, the $\mathsf{DP}\textsc{ReDI}_{1}$ set shifted substantially toward larger $C(1\to 2)$ and was the widest interval. This is because the DP projection with a higher $\beta$ is a more stable inferential target as seen in Figure 3, though this robustness may come at the expense of efficiency as also shown in pair 76.